Grade 1210 grade science

How can three resistors of resistances 2 Ω , 3 Ω , and 6 Ω be connected to give a total resistance of (a) 4 Ω

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student


When R2 and R3 are connected in parallel with R1 in series we get
1/R = 1/R2 + 1/R3
g= 1/3 + 1/6
= 1/2
Thus, R = 2 ohm
Resistance in series = R + R1
= 2 + 2
= 4 ohm

Thanks